solve the system by addition or substitution:
-15x+3y=6
y=2+5x

Answer by checkley71(8403)

y=2+5x
5x-y=-2 now multiply this equation by 3 & add it to the other equation.
15x-3y=-6
-15x+3y=6 now add
--------------------------------
0+0=0 these are identical equation therefore there is no unique solution.